if y=sqrt(x+1) + sqrt(x-1) prove that sqrt(x^2 -1)dy/dx=(1/2)y

Grade:12

2 Answers

Rinkoo Gupta
9 years ago

y=v(x+1) +v(x-1)

dy/dx=1/2v(x+1) +1/2v(x-1)

=(1/2)[(v(x-1)+v(x+1))}]/v(x-1)v(x+1)]

dy/dx.v(x^2-1)=y/2


Thanks & Regards

Rinkoo Gupta

AskIITians Faculty

Arun Kumar IIT Delhi
9 years ago
Hi


Thanks & Regards, Arun Kumar, Btech,IIT Delhi, Askiitians Faculty